OpenStudy (anonymous):

Let f(x)= sinx/x a.) What is the average rate of change of f on the interval [π/2, 3π/2]? b.) What is lim f(x) as x approaches 0? c.)Is the line x=0 a vertical asymptote of f? Justify your answer using limits. d.) Use the Squeeze Theorem to show the one y=0 is a horizontal asymptote of f.

OpenStudy (anonymous):

average rate of change of f(x) in the interval (a,b) is: f(b)-f(a)/b-a

OpenStudy (anonymous):

How do I know what f (π/2) is though?

OpenStudy (anonymous):

f(π/2) = sin π/2/π/2 = ?

OpenStudy (anonymous):

1/π/2

OpenStudy (anonymous):

Sorry, I'm missing my calculator right now -.-

OpenStudy (anonymous):

right: 2/π

OpenStudy (anonymous):

Ah, yes.

OpenStudy (anonymous):

do the same for the other point

OpenStudy (anonymous):

-2/3π?

OpenStudy (anonymous):

For sin(3π/2)/(3π/2)

OpenStudy (anonymous):

OH wait.... 3π/2 isn't even on the unit circle.

OpenStudy (anonymous):

it's the angle

OpenStudy (anonymous):

yeh it is, never mind...

OpenStudy (anonymous):

Yes, -2/3π

OpenStudy (anonymous):

f(3π/2) =-2/3π

OpenStudy (anonymous):

now do the ratio f(b)-f(a)/b-a

OpenStudy (anonymous):

Okay hold on.

OpenStudy (anonymous):

-7/2

OpenStudy (anonymous):

(-2/3π - 2/π)/(3π/2-π/2)=-8/3π/π=-8/3

OpenStudy (anonymous):

sigh... Okay :(

OpenStudy (anonymous):

The limit as X approaches 0 would be 1 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

X=0 is not a vertical asymptote is it?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

How can I justify that using limits?

OpenStudy (anonymous):

Because x approached infinity from both sides?

OpenStudy (anonymous):

infinity and negative infinity that is.

OpenStudy (anonymous):

no, lim f(x) for x->0 is 1, remmeber

OpenStudy (anonymous):

Yeh but how does that prove that x=0 is not a vertical asymptom?

OpenStudy (anonymous):

asymptote*

OpenStudy (anonymous):

when a lione is asymptot, the graph should stay in one of its sides, but in this case it is not true

OpenStudy (anonymous):

OKay!

OpenStudy (anonymous):

Can you give me a brief easy ideal of the "squeeze Theorem". i didn't really understand it in class.

OpenStudy (anonymous):

imagine 3 functions: f(x), g(x) and h(x) that have this property: f(x)< g(x) < h(x) when x->a limits of f(x) and g(x) are for example = b , so this will be also the limit of g(x)

OpenStudy (anonymous):

Hmmm... okay, so how can I use this Theorem to show that y=0 is a horizontal asymptote of f?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=f%28x%29%3D+sinx%2Fx i don't think y=0 is a horizontal asymptot

OpenStudy (anonymous):

So it would appear. Very interesting. Perhaps the worksheet has made an error. Regardless, thanks for your help!

OpenStudy (anonymous):

you wellcome